# Dusit Niyato, , Bo Ai

### v1

@@ III-A Centralized Processing
$\mathbf{v}_{k}=\mathbf{W}_{k}\hat{\mathbf{g}}_{k},$ (7)
where $\mathbf{W}_{k}\in\mathbb{C}^{MN\times MN}$ is an arbitrary matrix determined only by channel statistics instead of channel estimates. Alternatively, the conventional C-MMSE combining as $\mathbf{v}_{k}=p_{k}(\sum_{l=1}^{K}{p_{l}\left(\hat{\mathbf{g}}_{l}\hat{\mathbf{g}}_{l}^{H}+\mathbf{C}_{l}\right)}+\sigma^{2}\mathbf{I}_{MN})^{-1}\hat{\mathbf{g}}_{k}$ [13] is also on the form of (7), except that $\mathbf{W}_{k}$ depends on the instantaneous channel estimates. Note that $\mathbf{W}_{k}$ remains constant in a long period of time since the channel statistics change slowly in time [5]. Thus, by applying the combining scheme as in (7), only one matrix-vector multiplication is required for the combining scheme design for each UE in a coherence block.
$\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF}}=\frac{p_{k}\left|\mathrm{tr}\left(\mathbf{W}_{k}^{H}\bar{\mathbf{R}}_{k}\right)\right|^{2}}{\sum_{l=1}^{K}{p_{l}\mu_{kl}}+\sum_{l\in\mathcal{P}_{k}}{p_{l}\varepsilon_{kl}}+p_{k}\omega_{k}-p_{k}\left|\mathrm{tr}\left(\mathbf{W}_{k}^{H}\bar{\mathbf{R}}_{k}\right)\right|^{2}+\sigma^{2}\mathrm{tr}\left(\mathbf{W}_{k}\bar{\mathbf{R}}_{k}\mathbf{W}_{k}^{H}\right)}$ (8)
$\omega_{k}=\!p_{k}\tau_{p}\mathrm{tr(}\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})\mathrm{tr(}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})+p_{k}\tau_{p}\mathrm{tr(}\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{kk})\mathrm{tr(}\mathbf{W}_{k}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})-\!\mathrm{tr(}\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{kk}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})$ (10)
$+\sum\nolimits_{m_{1}=1}^{M}{\sum\nolimits_{m_{2}=1,m_{1}\neq m_{2}}^{M}{[\mathrm{tr}(\mathbf{W}_{k,m_{1}m_{1}}^{H}\bar{\mathbf{G}}_{m_{1}m_{2}kk}\mathbf{W}_{k,m_{2}m_{2}}\bar{\mathbf{G}}_{m_{1}m_{2}kk}^{H})}}+\mathrm{tr}(\mathbf{W}_{k,m_{2}m_{1}}^{H}\bar{\mathbf{G}}_{m_{2}kk}\mathbf{W}_{k,m_{2}m_{1}}\bar{\mathbf{G}}_{m_{1}kk})]$
$+\sum\nolimits_{m_{1}=1}^{M}{\mathrm{tr}(\mathbf{W}_{k,m_{1}m_{1}}^{H}\bar{\mathbf{G}}_{m_{1}kk}\mathbf{W}_{k,m_{1}m_{1}}\bar{\mathbf{G}}_{m_{1}kk})}$
$\mathbf{w}_{k}^{*}\!\!=\!\!\left(\sum_{l=1}^{K}{p_{l}\mathbb{E}\{(\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H})^{T}\otimes(\mathbf{g}_{l}\mathbf{g}_{l}^{H})\}}-p_{k}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})^{H}+(\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\}^{T}\otimes\mathbf{I}_{MN})\right)^{-1}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})$ (24)
When applying the combining scheme as in (7), we can compute (5) in novel closed-form as follows.
@@
With the combining scheme as in (7), the achievable SE for UE $k$ in (5) can be computed in closed-form as $\overline{\mathrm{SE}}_{k}^{\mathrm{c},\mathrm{UatF}}=\frac{\tau_{c}-\tau_{p}}{\tau_{c}}\log_{2}(1+\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF}})$ with $\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF}}$ given in (8) on the top of the next page, where $\varepsilon_{kl}=p_{k}p_{l}\tau_{p}^{2}|\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})|^{2}$,
$\mu_{kl}$ $=\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{ll}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})+\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\bar{\mathbf{G}}_{kk})$ (9)
$+\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{G}}_{ll}\mathbf{W}_{k}\hat{\mathbf{R}}_{k})+\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\hat{\mathbf{R}}_{k}),$
$\omega_{k}$ is given in (10) on the top of the next page, $\bar{\mathbf{G}}_{kk}=\mathrm{diag}[\bar{\mathbf{G}}_{1kk},\dots,\bar{\mathbf{G}}_{Mkk}]\in\mathbb{C}^{MN\times MN}$ with $\bar{\mathbf{G}}_{mkk}=\bar{\mathbf{g}}_{mk}\bar{\mathbf{g}}_{mk}^{H}\in\mathbb{C}^{N\times N}$, $\bar{\mathbf{R}}_{k}=\bar{\mathbf{G}}_{kk}+p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}\in\mathbb{C}^{MN\times MN}$, $\bar{\mathbf{G}}_{m_{1}m_{2}kk}=\bar{\mathbf{g}}_{m_{1}k}\bar{\mathbf{g}}_{m_{2}k}^{H}$, and $\mathbf{W}_{k,m_{1}m_{2}}\in\mathbb{C}^{N\times N}$ denotes the $(m_{1},m_{2})$-th sub-matrix of $\mathbf{W}_{k}$, respectively.
@@
Note that closed-form results in Theorem 1 are applicable for any deterministic $\mathbf{W}_{k}$. When $\mathbf{W}_{k}=\mathbf{I}_{MN}$, the results in Theorem 1 simplify to those of centralized MR (C-MR) combining such as [28, Corollary 2]. Moreover, the effects of the semi-deterministic LoS component can be clearly observed from (8), which are indicated by the $\bar{\mathbf{G}}_{kk}$-related terms. By letting $\bar{\mathbf{G}}_{kk}=\mathbf{0}$, we can derive the closed-form results for the Rayleigh fading channel model, which consists of only the NLoS component.
@@ III-B Distributed Processing
$\hat{x}_{mk}\!=\!\mathbf{v}_{mk}^{H}\mathbf{y}_{m}\!=\!\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}x_{k}\!\!+\!\!\sum_{l\neq k}^{K}{\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}x_{l}}\!+\!\mathbf{v}_{mk}^{H}\mathbf{n}_{m}.$ (11)
@@
$\mathbf{v}_{mk}=p_{k}\left(\sum_{l=1}^{K}{p_{l}\left(\hat{\mathbf{g}}_{ml}\hat{\mathbf{g}}_{ml}^{H}+\mathbf{C}_{ml}\right)}+\sigma^{2}\mathbf{I}_{N}\right)^{-1}\hat{\mathbf{g}}_{mk},$ (12)
which can minimize the local conditional MSE, $\mathrm{MSE}_{mk}=\mathbb{E}\{|x_{k}-\hat{x}_{mk}|^{2}|\hat{\mathbf{g}}_{mk}\}$. In the distributed processing scheme, the L-MMSE combining is widely implemented at each AP due to its outstanding performance [13, 27]. However, the L-MMSE combining is designed with $N\times N$ matrix inversion at every coherence block, which is computationally demanding when each AP is equipped with a large number of antennas, which might happen in 6G networks with high frequencies. Inspired by the structure of (12), we study a generalized distributed combining scheme like the form of (7) as
$\mathbf{v}_{mk}=\mathbf{W}_{mk}\hat{\mathbf{g}}_{mk},$ (13)
@@
After computing local data estimates as (11), all APs send them to the CPU to derive the final decoding by applying the LSFD weighting coefficients for UE $k$$\{a_{mk}:m=1,\dots,M\}$ estimates as $\check{x}_{k}=\sum_{m=1}^{M}{a_{mk}^{*}\hat{x}_{mk}}=(\sum_{m=1}^{M}{a_{mk}^{*}\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}})x_{k}+\sum_{l\neq k}^{K}{\sum_{m=1}^{M}{a_{mk}^{*}\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}x_{l}}}+\mathbf{n}_{k}^{\prime}$, where $\mathbf{n}_{k}^{\prime}=\sum_{m=1}^{M}{a_{mk}^{*}\mathbf{v}_{mk}^{H}\mathbf{n}_{m}}$. Note that only the decoding data estimates are sent to the CPU thus it does not know the instantaneous channel estimates. Thus, based on $\check{x}_{k}$, by applying the UatF capacity bound, we can derive the achievable SE expression for the distributed processing scheme as [13]
@@
where $\overline{\mathbf{b}}_{kk}=[\mathrm{tr}(\mathbf{W}_{1k}^{H}\bar{\mathbf{R}}_{1k}),\dots,\mathrm{tr}(\mathbf{W}_{Mk}^{H}\bar{\mathbf{R}}_{Mk})]^{T}\in\mathbb{C}^{M}$, and $\overline{\mathbf{D}}_{k}=\mathrm{diag}[\mathrm{tr}(\mathbf{W}_{1k}\bar{\mathbf{R}}_{1k}\mathbf{W}_{1k}^{H}),\dots,\mathrm{tr}(\mathbf{W}_{Mk}\bar{\mathbf{R}}_{Mk}\mathbf{W}_{Mk}^{H})]\in\mathbb{C}^{M\times M}$, respectively. As for $\overline{\mathbf{\Xi}}_{kl}\in\mathbb{C}^{M\times M}$, if $l\notin\mathcal{P}_{k}$, the $(n,m)$-th element of $\overline{\mathbf{\Xi}}_{kl}$ is $[\overline{\mathbf{\Xi}}_{kl}]_{nm}=\varphi_{mkl}$ for $m=n$ with
$\varphi_{mkl}$ $=\!\!\mathrm{tr}(\mathbf{W}_{mk}^{H}\bar{\mathbf{G}}_{mll}\mathbf{W}_{mk}\bar{\mathbf{G}}_{mkk})\!+\!\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{ml}\mathbf{W}_{mk}\bar{\mathbf{G}}_{mkk})$ (19)
$\!+\!\mathrm{tr}(\mathbf{W}_{mk}^{H}\bar{\mathbf{G}}_{mll}\mathbf{W}_{mk}\hat{\mathbf{R}}_{mk})+\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{ml}\mathbf{W}_{mk}\hat{\mathbf{R}}_{mk}),$
@@
$\begin{cases}0,\,\,\mathrm{if}\,\,l\in\mathcal{P}_{k}\backslash\left\{k\right\}\\ p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{mk}^{H}\bar{\mathbf{G}}_{mkk})\mathrm{tr}(\mathbf{W}_{mk}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk})\\ +p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk})\mathrm{tr}(\mathbf{W}_{mk}\bar{\mathbf{G}}_{mkk}),\,\,\mathrm{if}\,\,l=k\\ \end{cases}$
and $[\overline{\mathbf{\Xi}}_{kl}]_{nm}=\mathrm{tr}(\mathbf{W}_{mk}\mathbf{B}_{mkl})\mathrm{tr}(\mathbf{W}_{nk}^{H}\mathbf{B}_{nlk})$ for $m\neq n$ with $\mathbf{B}_{mlk}=\sqrt{p_{k}p_{l}}\tau_{p}\check{\mathbf{R}}_{ml}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk}$ for $l\in\mathcal{P}_{k}\backslash\{k\}$ and $\mathbf{B}_{mkk}=\bar{\mathbf{G}}_{mkk}+\sqrt{p_{k}p_{l}}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{ml}$ for $l=k$, and $\bar{\mathbf{R}}_{mk}=\bar{\mathbf{G}}_{mkk}+p_{k}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk}$, respectively.
@@
The derived closed-form results in Theorem 2 are applicable for any configuration of $\mathbf{W}_{mk}$. When $\mathbf{W}_{mk}=\mathbf{I}_{N}$, derived results can reduce to those of L-MR combining such as [26, Sec. III-A] for the L-MR combining based LSFD processing scheme over the Rician fading channel.

### v2

@@ III-A Centralized Processing
$\mathbf{v}_{k}=\mathbf{W}_{k}\widehat{\mathbf{g}}_{k},$ (7)
where $\mathbf{W}_{k}\in\mathbb{C}^{MN\times MN}$ is an arbitrary matrix determined only by channel statistics instead of channel estimates. Alternatively, the conventional C-MMSE combining as $\mathbf{v}_{k}=p_{k}(\sum_{l=1}^{K}{p_{l}\left(\widehat{\mathbf{g}}_{l}\widehat{\mathbf{g}}_{l}^{H}+\mathbf{C}_{l}\right)}+\sigma^{2}\mathbf{I}_{MN})^{-1}\widehat{\mathbf{g}}_{k}$ [16] is also on the form of (7), except that $\mathbf{W}_{k}$ in the C-MMSE combining depends on the instantaneous channel estimates. Note that $\mathbf{W}_{k}$ remains constant in a long period of time since the channel statistics change slowly in time [5]. Thus, by applying the combining scheme in (7), only one matrix-vector multiplication is required for the combining scheme design for each UE in a coherence block, avoiding computationally demanding instantaneous information-based matrix inversion.
When applying the combining scheme in (7), we can compute (5) in novel closed-form as follows.
@@
With the combining scheme in (7), the achievable SE for UE $k$ in (5) can be computed in closed-form as $\overline{\mathrm{SE}}_{k}^{\mathrm{c},\mathrm{UatF}}=\frac{\tau_{c}-\tau_{p}}{\tau_{c}}\log_{2}(1+\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF}})$ with $\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF}}$ given in (8) on the top of this page, where $\varepsilon_{kl}=p_{k}p_{l}\tau_{p}^{2}|\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k})|^{2}$,
$\mu_{kl}$ $=\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{ll}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})+\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\overline{\mathbf{G}}_{kk})$ (9)
$+\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{G}}_{ll}\mathbf{W}_{k}\widehat{\mathbf{R}}_{k})+\mathrm{tr}(\mathbf{W}_{k}^{H}\check{\mathbf{R}}_{l}\mathbf{W}_{k}\widehat{\mathbf{R}}_{k}),$
$\omega_{k}$ is given in (10) on the top of this page, $\overline{\mathbf{G}}_{kk}=\mathrm{diag}[\overline{\mathbf{G}}_{1kk},\dots,\overline{\mathbf{G}}_{Mkk}]\in\mathbb{C}^{MN\times MN}$ with $\overline{\mathbf{G}}_{mkk}=\overline{\mathbf{g}}_{mk}\overline{\mathbf{g}}_{mk}^{H}\in\mathbb{C}^{N\times N}$, $\overline{\mathbf{R}}_{k}=\overline{\mathbf{G}}_{kk}+p_{k}\tau_{p}\check{\mathbf{R}}_{k}\mathbf{\Psi}_{k}^{-1}\check{\mathbf{R}}_{k}\in\mathbb{C}^{MN\times MN}$, $\overline{\mathbf{G}}_{m_{1}m_{2}kk}=\overline{\mathbf{g}}_{m_{1}k}\overline{\mathbf{g}}_{m_{2}k}^{H}$, and $\mathbf{W}_{k,m_{1}m_{2}}\in\mathbb{C}^{N\times N}$ denotes the $(m_{1},m_{2})$-th sub-matrix of $\mathbf{W}_{k}$, respectively.
@@
Note that closed-form results in Theorem 1 are applicable for any deterministic $\mathbf{W}_{k}$. When $\mathbf{W}_{k}=\mathbf{I}_{MN}$, the results in Theorem 1 simplify to those of centralized MR (C-MR) combining such as [32, Corollary 2]. Moreover, the effects of the semi-deterministic LoS component can be clearly observed from (8), which are indicated by the $\overline{\mathbf{G}}_{kk}$-related terms. By letting $\overline{\mathbf{G}}_{kk}=\mathbf{0}$, we can derive the closed-form results for the Rayleigh fading channel model, which consists of only the NLoS component.
@@ III-B Distributed Processing
$\widehat{x}_{mk}\!=\!\mathbf{v}_{mk}^{H}\mathbf{y}_{m}\!=\!\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}x_{k}\!\!+\!\!\sum_{l\neq k}^{K}{\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}x_{l}}\!+\!\mathbf{v}_{mk}^{H}\mathbf{n}_{m}.$ (11)
@@
$\mathbf{v}_{mk}=p_{k}\left(\sum_{l=1}^{K}{p_{l}\left(\widehat{\mathbf{g}}_{ml}\widehat{\mathbf{g}}_{ml}^{H}+\mathbf{C}_{ml}\right)}+\sigma^{2}\mathbf{I}_{N}\right)^{-1}\widehat{\mathbf{g}}_{mk},$ (12)
which can minimize the local conditional MSE, $\mathrm{MSE}_{mk}=\mathbb{E}\{|x_{k}-\widehat{x}_{mk}|^{2}|\widehat{\mathbf{g}}_{mk}\}$. In the distributed processing scheme, the L-MMSE combining is widely implemented at each AP due to its outstanding performance [16, 30]. However, the L-MMSE combining is designed with $N\times N$ matrix inversion at every coherence block, which is computationally demanding when each AP is equipped with a large number of antennas, which might happen in 6G networks with high frequencies. Inspired by the structure of (12), we study a generalized distributed combining scheme like the form of (7) as
$\mathbf{v}_{mk}=\mathbf{W}_{mk}\widehat{\mathbf{g}}_{mk},$ (13)
@@
After computing local data estimates as (11), all APs send them to the CPU to derive the final decoding by applying the LSFD weighting coefficients for UE $k$$\{a_{mk}:m=1,\dots,M\}$ estimates as $\check{x}_{k}=\sum_{m=1}^{M}{a_{mk}^{*}\widehat{x}_{mk}}=(\sum_{m=1}^{M}{a_{mk}^{*}\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}})x_{k}+\sum_{l\neq k}^{K}{\sum_{m=1}^{M}{a_{mk}^{*}\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}x_{l}}}+\mathbf{n}_{k}^{\prime}$, where $\mathbf{n}_{k}^{\prime}=\sum_{m=1}^{M}{a_{mk}^{*}\mathbf{v}_{mk}^{H}\mathbf{n}_{m}}$. Note that only the decoding data estimates are sent to the CPU thus it does not know the instantaneous channel estimates. Thus, based on $\check{x}_{k}$, by applying the UatF capacity bound, we can derive the achievable SE expression for the distributed processing scheme as [16]
@@
where $\overline{\mathbf{b}}_{kk}=[\mathrm{tr}(\mathbf{W}_{1k}^{H}\overline{\mathbf{R}}_{1k}),\dots,\mathrm{tr}(\mathbf{W}_{Mk}^{H}\overline{\mathbf{R}}_{Mk})]^{T}\in\mathbb{C}^{M}$ and $\overline{\mathbf{D}}_{k}=\mathrm{diag}[\mathrm{tr}(\mathbf{W}_{1k}\overline{\mathbf{R}}_{1k}\mathbf{W}_{1k}^{H}),\dots,\mathrm{tr}(\mathbf{W}_{Mk}\overline{\mathbf{R}}_{Mk}\mathbf{W}_{Mk}^{H})]$ $\in\mathbb{C}^{M\times M}$, respectively. As for $\overline{\mathbf{\Xi}}_{kl}\in\mathbb{C}^{M\times M}$, if $l\notin\mathcal{P}_{k}$, the $(n,m)$-th element of $\overline{\mathbf{\Xi}}_{kl}$ is $[\overline{\mathbf{\Xi}}_{kl}]_{nm}=\varphi_{mkl}$ for $m=n$ with
$\varphi_{mkl}$ $=\!\!\mathrm{tr}(\mathbf{W}_{mk}^{H}\overline{\mathbf{G}}_{mll}\mathbf{W}_{mk}\overline{\mathbf{G}}_{mkk})\!+\!\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{ml}\mathbf{W}_{mk}\overline{\mathbf{G}}_{mkk})$ (19)
$\!+\!\mathrm{tr}(\mathbf{W}_{mk}^{H}\overline{\mathbf{G}}_{mll}\mathbf{W}_{mk}\widehat{\mathbf{R}}_{mk})+\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{ml}\mathbf{W}_{mk}\widehat{\mathbf{R}}_{mk}),$
@@
$\begin{cases}0,\,\,\mathrm{if}\,\,l\in\mathcal{P}_{k}\backslash\left\{k\right\}\\ p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{mk}^{H}\overline{\mathbf{G}}_{mkk})\mathrm{tr}(\mathbf{W}_{mk}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk})\\ +p_{k}\tau_{p}\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk})\mathrm{tr}(\mathbf{W}_{mk}\overline{\mathbf{G}}_{mkk}),\,\,\mathrm{if}\,\,l=k\\ \end{cases}$
and $[\overline{\mathbf{\Xi}}_{kl}]_{nm}=\mathrm{tr}(\mathbf{W}_{mk}\mathbf{B}_{mkl})\mathrm{tr}(\mathbf{W}_{nk}^{H}\mathbf{B}_{nlk})$ for $m\neq n$ with $\mathbf{B}_{mlk}=\sqrt{p_{k}p_{l}}\tau_{p}\check{\mathbf{R}}_{ml}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk}$ for $l\in\mathcal{P}_{k}\backslash\{k\}$ and $\mathbf{B}_{mkk}=\overline{\mathbf{G}}_{mkk}+\sqrt{p_{k}p_{l}}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{ml}$ for $l=k$, and $\overline{\mathbf{R}}_{mk}=\overline{\mathbf{G}}_{mkk}+p_{k}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk}$, respectively.
@@
Note that (15) is derived based on the UatF capacity bound, which holds for arbitrary receive combining vectors $\mathbf{v}_{mk}$. Indeed, for each realization of the AP/UE locations, (15) can be computed via the Monte-Carlo simulation methodology by considering many instantaneous channel realizations. When the OBE combining scheme $\mathbf{v}_{mk}=\mathbf{W}_{mk}\widehat{\mathbf{g}}_{mk}$ in (13) is applied, the expectations in (15) can be computed in closed-form in (18) based on the channel statistics.
Remark 4.
The derived closed-form results in Theorem 2 are applicable for any configuration of $\mathbf{W}_{mk}$. When $\mathbf{W}_{mk}=\mathbf{I}_{N}$, derived results can reduce to those of L-MR combining such as [29, Sec. III-A] for the L-MR combining based LSFD processing scheme over the Rician fading channel.